OpenStudy (anonymous):

13/12 is this a reduced fraction?

OpenStudy (dmb1215):

no its would be 1 1/12

OpenStudy (anonymous):

12 goes into 13 one time and you keep the denominator so it is keep same denominator. 1 1/12

OpenStudy (anonymous):

The male-to-female ratio in Taiwan is 52 to 48. Write this ratio as a fraction in reduced form. 52/48 52/4=13 48/2=12 12/13=1r1 then I reduce it to a mixed number 1 1/12, I have to show my work does this look correct

OpenStudy (dmb1215):

yep

OpenStudy (anonymous):

thanks :)

OpenStudy (triciaal):

should be 13/12

OpenStudy (triciaal):

leave as a reduced fraction